OpenStudy (anonymous):

Find the values of a and b that make f continuous everywhere. f(x) = (x^2-9)/(x-3) if x<3 ax^2-bx+3 if 3<=x<4 2x-a+b if x>=4

OpenStudy (anonymous):

the two answers i got were, b=-1, a=0. what i did was set: (x^2-9)/(x-3)=ax^2-bx+3 and from this i got b=3a-1, and then i set ax^2-bx+3=2x-a+b.

OpenStudy (anonymous):

find limit left hand limit of x->3 lim x->3 (x+3(x-3)/(x-3)=6 if it has to be continuous then left and right limits must be equal . right hand limits of x->3 a(3)^2-3b+3=6 9a-3b=3 now left and right limits of x->4 left hand limit 16a-4b+3 right hand limit 8-a+b so they should be equal 16a-4b+3=8-a+b 17a-5b=5 solve the following two equations 9a-3b=3 17a-5b=5 which results the same a=0 b=-1 you got it correct !

OpenStudy (anonymous):

okay tyvm